I am having a really difficult time with this problem - it is not in a book it is homework and I am lost.
The probability that a life insurance salesperson following up a magazine lead will make a sale is 30%. A salesperson has two leads on a certain day. What is the probability that the salesperson will sell (i) both? (ii) exactly one policy; (iii) at least one policy?

Answer by stanbon(48568)

The probability that a life insurance salesperson following up a magazine lead will make a sale is 30%.
P(sale)= 0.3 ; P(no sale) = 0.7
-------------------------
A salesperson has two leads on a certain day. What is the probability that the salesperson will sell (i) both?
(.3)^2 = = 0.09
-------------------------
(ii) exactly one policy;
P(exactly one ) = 2C1(0.3)*(0.7)= 2*0.21 = 0.42
-------------------------
(iii) at least one policy?
P(at least one sale) = 1 - P(no sales) = 1-(0.7)^2 = 0.51
-------------
Cheers,
Stan H.